OpenStudy (anonymous):

If x/y=5/8, then x/5=

OpenStudy (anonymous):

If x/5 = x/8, then x/5 - x/8 = 0. x/5 - x/8 = 0 (8x)/40 - (5x)/40 = 0 (3x)/40 = 0 3x = 0 x = 0 So (x+5):(x+8) = (0+5):(0+8) = 5:8.